# On the Generalization Error of Differentially Private Algorithms Via Typicality

### v1

@@ III Mutual Information Bounds
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$. Fix $|\mathcal{Z}|\geq 2$ and $N\geq 1$. For every $t\in\{1,2,\dots,\lfloor 2\sqrt{N\log N}\rfloor\}$ and every dataset $s\in\mathcal{S}$, there exists a distribution $Q_{W}^{(t)}$ such that
$\quad\,\,\,\,\,\,+\frac{2|\mathcal{Z}|\varepsilon}{N}+\mathds{1}_{\{t\geq 2\}}\min\{|\mathcal{Z}|\log t,\log(|\mathcal{Z}|t^{|\mathcal{Z}|-1})\}$ (11)
where $\Delta_{t}:=\lceil 2\sqrt{N\log N}/t\rceil+1$.
@@
We plot two representative scenarios and show a comparison between our bounds and the bounds in [1].
@@
    Figure 1: a small dataset with $N=1\times 10^{3}$ and $|\mathcal{Z}|=2$. The setting is typical in applications with limited sample sizes and binary outcomes/features, such as clinical studies or small A/B tests, where we collect data from thousands of patients/users and each record carries a binary attribute (e.g., responder/non-responder, click/no-click).
@@
    Figure 2: a large dataset with $N=1\times 10^{7}$ and $|\mathcal{Z}|=1\times 10^{6}$. This regime arises in large-scale telemetry or interaction logs with high-cardinality categorical values, such as search queries, URLs, item IDs, or token IDs, where one observes tens of millions of events drawn from a vocabulary on the order of millions.
Figure 1: Dataset is of size $1\times 10^{3}$ with $|\mathcal{Z}|=2$.
Figure 2: Dataset is of size $1\times 10^{7}$ with $|\mathcal{Z}|=1\times 10^{6}$.
@@ IV Maximal Leakage Bounds
In this section we consider the bound on maximal leakage, also by a typicality-based approach. By (3), these upper bounds directly translate to generalization error bounds.
@@
Note $\mathcal{L}(S\to W)$ is the Sibson’s mutual information [48, 49] of order infinity, and it upper bounds mutual information. For an $\varepsilon$-DP mechanism, in nats, we have [15]
@@
The key advantage of maximal leakage is that it depends on the sample distribution only through its support. Consequently, it is independent of the precise distribution over the samples and is particularly amenable to analysis, especially in additive noise settings [18, 29].
Note that the usual way to keep the maximal leakage of an algorithm bounded is to add noise (e.g., Laplacian or Gaussian [18]), which aligns with common differential privacy mechanisms [12]. Hence, for the study of the generalization error of differentially private algorithms, compared to bounding mutual information, it is more natural to bound the maximal leakage.
We introduce a simple bound that will be useful.
@@ IV-A General Algorithms
We first provide a general bound on maximal leakage, where $\mathcal{A}=P_{W|S}$ is not necessarily differentially private.
@@
Note above result directly recovers [1, Proposition 1] (also see (6)), considering maximal leakage upper-bounds mutual information. Hence we provide an alternative proof.
If we utilize the typicality bound in our Claim 2 instead, above Theorem 8 can be strengthened to
@@ IV-B Differentially Private Algorithms
We then consider $\varepsilon$-differentially private algorithms. In the regime of $\varepsilon\leq 1$, we derive bounds that are tighter than the general Theorem 8; otherwise, Theorem 8 is better. This phenomenon has also been observed in [1].
@@
Let $S=(Z_{1},\ldots,Z_{N})$ be a dataset of size $N$ where $Z_{i}\in\mathcal{Z}$ are sampled i.i.d. from $P_{Z}$. Let $W\in\mathcal{W}$ be a hypothesis obtained from an algorithm $\mathcal{A}$ whose induced distribution is $P_{W|S}$. For all $s\in\mathcal{S}$, if $\mathcal{A}$ is $\varepsilon$-DP,
@@
    $\mathcal{L}(S\rightarrow W)\leq(|\mathcal{Z}|-1)N\varepsilon.$
@@
The proof of Theorem 9 can be found in Appendix H. The proof relies on a special choice of representative sets. This approach can be combined with the refined constructions of representative sets developed in the proofs of Theorems 5 and 6. Consequently, it is possible to integrate these constructions into the current proof of Theorem 9 and further strengthen our maximal leakage bounds. We leave this direction for future work.
@@ V Concluding Remarks
We have provided two classes of explicit generalization error bounds, one based on mutual information and another based on maximal leakage. For the former class, we strictly improve the bounds in [1], while for the latter class, we propose novel bounds. It is nontrivial but of interest to apply the techniques in Section III to the maximal leakage bounds in Section IV. It is possible to improve bounds in both classes through better choices of the mixture distribution in Lemma 2.
Moreover, in [1], another class of private algorithms based on Gaussian differential privacy [44] has been studied. We will also study them in the future, together with other extensions of differential privacy [42, 43].
@@ Appendix E Proof of Theorem 6
$P_{Z}^{\otimes N}(\mathcal{Z}\backslash\mathcal{T})\leq 2|\mathcal{Z}|\exp(-2N\varepsilon^{2}).$ (22)
@@
$I(S;W)$ $\leq\mathbb{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\in\mathcal{T}\right]\mathbb{P}(S\in\mathcal{T})$
$\,\,\,\,\,+\mathbb{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\notin\mathcal{T}\right]\mathbb{P}(S\notin\mathcal{T}).$ (23)
@@
$\mathbb{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\notin\mathcal{T}\right]\mathbb{P}(S\notin\mathcal{T})\leq\frac{2m\varepsilon}{N}.$

### v2

@@ III Mutual Information Bounds
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$. Fix $|\mathcal{Z}|\geq 2$ and $N\geq 1$. For every dataset $s\in\mathcal{S}$, there exists a distribution $Q_{W}^{(t)}$ such that
$I(S;W)\leq\min_{t\in\{1,2,\dots,\lfloor 2\sqrt{N\log N}\rfloor\}}\Big\{\frac{\varepsilon|\mathcal{Z}|\sqrt{N\log N}}{t}$
$\quad\quad\quad\quad-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon(\lceil 2\sqrt{N\log N}/t\rceil+1)}\bigr)+\frac{2|\mathcal{Z}|\varepsilon}{N}$
$\quad\quad\quad\quad+\mathds{1}_{\{t\geq 2\}}\min\{|\mathcal{Z}|\log t,\log(|\mathcal{Z}|t^{|\mathcal{Z}|-1})\}\Big\}$ (12)
@@
We plot two representative scenarios and show comparison between our bounds and the bounds in [1].
@@
    Figure 1: A small dataset with $N=1\times 10^{3}$ and $|\mathcal{Z}|=2$. The setting is typical in applications with limited sample sizes and binary outcomes/features, such as clinical studies or small A/B tests, where we collect data from thousands of patients/users and each record carries a binary attribute (e.g., responder/non-responder, click/no-click). Figure 1 shows that Theorem 4, 5 and 6 strictly improve Proposition $2$, 33 and $4$ in [1], respectively, and our Theorem 6 provides the tightest guarantee.
@@
    Figure 2: A large dataset with $N=1\times 10^{7}$ and $|\mathcal{Z}|=1\times 10^{6}$. This regime arises in large-scale telemetry or interaction logs with high-cardinality categorical values, such as search queries, URLs, item IDs, or token IDs, where one observes tens of millions of events drawn from a vocabulary on the order of millions. Figure 2 shows that in this case we only substantially improve the bounds in [1] when the privacy budget $\epsilon$ is small.
Figure 1: Dataset of size $N=1\times 10^{3}$ with $|\mathcal{Z}|=2$.
Figure 2: Dataset of size $N=1\times 10^{7}$ with $|\mathcal{Z}|=1\times 10^{6}$.
@@ IV Maximal Leakage Bounds
In this section, we present novel generalization bounds that can be efficiently computed by bounding maximal leakage and utilizing(3), also based on a typicality-based approach.
@@
For discussions on maximal leakage for continuous channels, see [15]. Note $\mathcal{L}(S\to W)$ is the Sibson’s mutual information [47, 48] of order infinity, and it upper bounds mutual information. For an $\varepsilon$-DP mechanism, we have [15]
@@
The key advantage of maximal leakage is that it depends on the sample distribution only through its support. Consequently, the corresponding generalization error bounds are independent of the precise distribution over the samples and are particularly amenable to analysis, especially in additive noise settings. For generalization bounds via maximal leakage, see [18, 22]
Moreover, note that the usual way to keep the maximal leakage of an algorithm bounded is to add noise (e.g., Laplacian or Gaussian [18]), which aligns with common differential privacy mechanisms [13]. Hence, for the study of the generalization error of differentially private algorithms, compared to bounding mutual information, it is more natural to bound the maximal leakage.
We then introduce a simple bound that will be useful.
@@ IV-A General Algorithms
We first provide a general upper bound on maximal leakage, where $\mathcal{A}=P_{W|S}$ is not necessarily differentially private.
@@
Theorem 8 exactly recovers [1, Proposition 1] (also see (6)) since maximal leakage upper-bounds mutual information strictly.
If we utilize the typicality bound in our Claim 2 instead, Theorem 8 can be strengthened to
@@ IV-B Differentially Private Algorithms
We then consider $\varepsilon$-differentially private algorithms. In the regime of $\varepsilon\leq 1$, we derive the following results that are tighter than Theorem 8.
@@
Let $S=(Z_{1},\ldots,Z_{N})$ be a dataset of size $N$ where $Z_{i}\in\mathcal{Z}$ are sampled i.i.d. from $P_{Z}$. Let $W\in\mathcal{W}$ be a hypothesis obtained from an $\varepsilon$-DP algorithm $\mathcal{A}$ whose induced distribution is $P_{W|S}$. For all datasets $s\in\mathcal{S}$,
@@
    $\mathcal{L}(S\rightarrow W)\leq(|\mathcal{Z}|-1)N\varepsilon,$
@@
The proof of Theorem 9 can be found in Appendix H.
The proof relies on a special choice of representative sets. This approach can be combined with the refined constructions of representative sets developed in the proofs of Theorems 5 and 6. Consequently, it is possible to integrate these constructions for mutual information bounds into the current proof of Theorem 9 and further strengthen our maximal leakage bounds. We leave this direction for future work.
@@ V Concluding Remarks
In this paper, we have provided two classes of explicit, easily computable generalization error bounds: one based on mutual information and the other based on maximal leakage. For the former class, we strictly improved the bounds in [1], and the comparison is shown in Figure 1 and Figure 2. Application of the techniques in Section III to the maximal leakage bounds in Section IV is non-trivial and left for future work. Moreover, it is still possible to improve bounds in both classes through better choices of the mixture distribution in Lemma 2.
In [1], another class of private algorithms based on Gaussian differential privacy [44] has been studied. It is also of interest to study the generalization error of Gaussian DP algorithms as well, together with other extensions of differential privacy [42, 43]. If specific privatization mechanisms are specified, sharper bounds are expected to be obtainable.
@@ Appendix E Proof of Theorem 6
$P_{Z}^{\otimes N}(\mathcal{Z}^{N}\backslash\mathcal{T})\leq 2|\mathcal{Z}|\exp(-2N\varepsilon^{2}).$ (23)
@@
$I(S;W)$ $\leq\mathbf{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\in\mathcal{T}\right]\mathbb{P}(S\in\mathcal{T})$
$\,\,\,\,\,+\mathbf{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\notin\mathcal{T}\right]\mathbb{P}(S\notin\mathcal{T}).$ (24)
@@
$\mathbf{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\notin\mathcal{T}\right]\mathbb{P}(S\notin\mathcal{T})\leq\frac{2m\varepsilon}{N}.$